# Language Models as Causal Effect Generators

### v1

@@ 3.2 Sequence-driven structural causal models (SD-SCMs)
Definition 6 (Parent-only concatenation)
Given DAG $\mathcal{G}$ over $m$ sequence variables $(\tilde{v}_{1},\ldots,\tilde{v}_{m})$ and a topological ordering $\tau$ consistent with $\mathcal{G}$, parent-only concatenation defines $\left(\tilde{v}_{i}\mid\mathbf{PA}_{i}\right)\equiv\bigoplus_{t\in\mathbf{PA}_{i}}\tilde{v}_{t}$, where $\mathbf{PA}_{i}$ are the parents of $\tilde{v}_{i}$ in $\mathcal{G}$ ordered according to $\tau$.
@@
Definition 7 (Sequence-driven structural causal model (SD-SCM))
@@
An SD-SCM uses $\mathcal{P}$ to define an observational distribution over the variables in $\mathbf{V}\cup\mathbf{U}$ that factorizes according to $\mathcal{G}$ via domain-restricted ancestral sampling and parent-only concatenation with $\tau$: $P^{\mathfrak{B}}\equiv\prod_{\tilde{x}_{t}\in\tau}\mathcal{P}_{\tilde{x}_{t}\mid C,\mathbf{PA}_{t}}$.
@@
The key difference between an SCM and an SD-SCM is that all variables have at least one common ancestor — the prior inputs $C$ that were used to train the language model. As with the observational distribution, domain-restricted ancestral sampling and parent-only concatenation also allow us to define interventional and counterfactual distributions for an SD-SCM.
Definition 8 (Sequence-driven interventional distribution)
@@
An SD-SCM $\mathfrak{B}$ entails the distribution of any subset of variables in $\mathbf{V}\cup\mathbf{U}$ following intervention $I=\text{do}\left(\tilde{v}_{i}=v\right)$ by replacing variable $\tilde{v}_{i}$ with value $v$, and otherwise sampling in the same manner. As with an SCM, we denote an SD-SCM after intervention $I$ as $\mathfrak{B}^{\text{do}(I)}$ and the resulting interventional distribution as $P^{\mathfrak{B};\text{do}(I)}$. This is computed for intervention $\text{do}(\tilde{v}_{i}=v)$ as follows: $P^{\mathfrak{B};\text{do}(\tilde{v}_{i}=v)}\equiv\prod_{\tilde{x}_{t}\in\tau}\mathcal{P}_{\tilde{x}_{t}\mid C,\tilde{v}_{i}=v,\mathbf{PA}^{\prime}_{t}}$, where $\mathbf{PA}^{\prime}_{t}$ denotes the parents of $\tilde{x}_{t}$, excluding $\tilde{v}_{i}$ if present.
@@
In order to admit unique answers to counterfactual queries, we define abduction for an SD-SCM given evidence $\mathbf{Z}=\mathbf{z}$ as the setting of values $\mathbf{U}=\mathbf{u}$ and any evidence upstream of the intervention, rather than a distribution $P_{\mathbf{U}\mid\mathbf{Z}=\mathbf{z}}$. [^4] Other choices can be explored here, which we leave to future extensions of SD-SCMs. In order to obtain such values $\mathbf{u}$, one needs access to more than just the observed data and language model $\mathcal{P}$ — obtaining $\mathbf{u}$ requires performing bookkeeping during the data generation process. This is a restatement of the fact that computing point counterfactuals in SCMs requires causal mechanisms that are invertible with respect to the noise variables in order to uniquely reconstruct the noise that produced a given observation. Because our primary application of SD-SCMs in this work is data generation, such bookkeeping is possible in all our use cases.
Definition 9 (Sequence-driven counterfactual distribution)
@@
Counterfactual sequence variable $\mathbf{Y}^{*}$ given factual evidence $\mathbf{z}$ and intervention $\text{do}(\mathbf{X}=\mathbf{x})$ (where $\mathbf{X},\mathbf{Y},\mathbf{Z}\subseteq\mathbf{V}$) can be computed for an SD-SCM $\mathfrak{B}$ whenever the exogenous setting $\mathbf{u}=\{u_{1},u_{2},\ldots,u_{|\mathbf{U}|}\}$ that generated evidence $\mathbf{z}$ is known. As with an SCM, for intervention $I$ and observed $\mathbf{Z}=\mathbf{z}$, we denote the counterfactual distribution $P^{\mathfrak{B}\mid\mathbf{Z}=\mathbf{z};\text{do}(I)}$. This is computed for evidence $\mathbf{z}$, exogenous conditions $\mathbf{u}$, and intervention $\text{do}(\tilde{v}_{i}=v)$ as follows: $P^{\mathfrak{B}\mid\mathbf{Z}=\mathbf{z};\text{do}(\tilde{v}_{i}=v)}\equiv\prod_{\tilde{x}_{t}\in\tau}\mathcal{P}_{\tilde{x}_{t}\mid C,\mathbf{U}=\mathbf{u},\mathbf{Z}^{\prime}=\mathbf{z}^{\prime},\tilde{v}_{i}=v,\mathbf{PA}^{\prime\prime}_{t}}$, where $\mathbf{Z}^{\prime}\subseteq\mathbf{Z}$ contains all ancestors of $\tilde{v}_{i}$ present in $\mathbf{Z}$, and $\mathbf{PA}^{\prime\prime}_{t}=\mathbf{PA}_{t}\setminus\left(\mathbf{U}\cup\mathbf{Z}^{\prime}\cup\{\tilde{v}_{i}\}\right)$.
@@
In summary, data can be generated from an SD-SCM by domain-restricted forward sampling variables in topological order, and, with adequate bookkeeping, both interventional and counterfactual samples can also be drawn. The key difficulty in generating data this way that is also useful for benchmarking causal inference methods is to ensure it has meaningful structure. In short, it is easy to generate data, but more difficult to generate useful data. The reason for this challenge is that we do not directly specify the structural equations; rather, the structural equations are determined by whatever the language model $\mathcal{P}$ has already encoded.
3.3 Example SD-SCM: going to the gym to improve your marathon time
Let sequence variables $\mathbf{U}=\{\tilde{u}_{1},\tilde{u_{2}}\}$ and $\mathbf{V}=\{\tilde{w}=$weather, $\tilde{g}=$gymOrRun, $\tilde{m}=$ marathon}}\}} be defined over the following sample spaces.
[…]
For this example, we use GPT-2 [49] as language model $\mathcal{P}$ and consider whether going for a run outside ($\tilde{g}=1$) will worsen marathon running times ($\tilde{m}=1$), compared to going to the gym ($\tilde{g}=0$). Figure 1 demonstrates two SD-SCMs ($\mathfrak{B}_{1}$ and $\mathfrak{B}_{2}$), corresponding to two different DAGs, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. $\mathcal{G}_{1}$ makes weather a confounder: $\tilde{u}_{1}\rightarrow\texttt{gymOrRun}\leftarrow\texttt{weather}\rightarrow\texttt{marathon}\leftarrow\tilde{u}_{2}.$ $\mathcal{G}_{2}$ instead makes weather a collider: $\tilde{u}_{1}\rightarrow\texttt{gymOrRun}\rightarrow\texttt{weather}\leftarrow\texttt{marathon}\leftarrow\tilde{u}_{2}.$ Therefore, $P^{\mathfrak{B}_{1}}$ factorizes according to $\mathcal{G}_{1}$ as $P^{\mathfrak{B}_{1}}=\mathcal{P}_{\tilde{g}\mid C,\tilde{w},\tilde{u}_{1}}\mathcal{P}_{\tilde{m}\mid C,\tilde{w},\tilde{u}_{2}}\mathcal{P}_{\tilde{w}\mid C}\mathcal{P}_{\tilde{u}_{1}\mid C}P_{\tilde{u}_{2}},$ where we use a uniform distribution for $P_{\tilde{u}_{2}}$, as shown in Figure 1(a). $P^{\mathfrak{B}_{2}}$ instead factorizes according to $\mathcal{G}_{2}$: $P^{\mathfrak{B}_{2}}=\mathcal{P}_{\tilde{w}\mid C,\tilde{g},\tilde{m}}\mathcal{P}_{\tilde{g}\mid C,\tilde{u}_{1}}\mathcal{P}_{\tilde{m}\mid C,\tilde{u}_{2}}\mathcal{P}_{\tilde{u}_{1}\mid C}P_{\tilde{u}_{2}}$. Under $P^{\mathfrak{B}_{1}}$, we would
Figure 2: Estimation error for random forest predictions of the SATE, with and without including weather $\tilde{w}$.
expect the magnitude of correlation $\rho$ between $Pr(\tilde{g}=1)$ (the propensity of running) and $Pr(\tilde{m}=1)$ (the propensity of marathon worsening) to decrease if we condition on confounder $\tilde{w}$. Under $P^{\mathfrak{B}_{2}}$, where $\tilde{w}$ is a collider, we would instead expect the magnitude of $\rho$ to increase. Figure 1(b) shows that this is indeed the case, computed across 1000 samples drawn from each observational distribution. We can see how this would, in turn, impact causal effect estimation by drawing the corresponding 1000 counterfactual samples from $P^{\mathfrak{B}\mid\mathbf{V}=\mathbf{v};\text{do}(\tilde{g}=0)}$ and/or $P^{\mathfrak{B}\mid\mathbf{V}=\mathbf{v};\text{do}(\tilde{g}=1)}$ for each observed unit and comparing estimates of, e.g., the sample average treatment effect (SATE) to the true SATE. The SATE describes the average effect over the given sample, which we can calculate directly as the observed units’ average ITE. Letting $\tilde{y}_{i}(1),\tilde{y}_{i}(0)$ denote the outcomes of the $i$-th unit given $\text{do}(\tilde{g}=1)$ and $\text{do}(\tilde{g}=0)$, respectively, we calculate the SATE across $n$ units as $\frac{1}{n}\sum_{i=1}^{n}\left(\tilde{y}_{i}(1)-\tilde{y}_{i}(0)\right)$. Figure 2 shows prediction error (in standard deviation units) when using a random forest to predict the SATE with $P(\texttt{marathon}=1)$ as the outcome, either using treatment $\tilde{g}$ as the only covariate, or using both $\tilde{g}$ and $\tilde{w}$. As we would expect, including a confounder leads to more accurate effect estimation, while including a collider does not.
4 Generating a benchmark for causal effect estimation
Figure 3: (a) A useful DAG structure for an SD-SCM-generated estimation benchmark. (b) Visual depiction of the structure in (a) used to create the breast cancer SD-SCM in Section 4.1.
In approaching the design of an SD-SCM-generated benchmark, we focus on the fully sequential DAG structure shown in Figure LABEL:fig:benchmark_dag. Exogenous variables $\mathbf{U}$ precede covariates $\mathbf{X}$, which in turn precede treatment $\tilde{t}$. All variables precede outcome $\tilde{y}$. Recall that the presence of an edge in a DAG allows for the possibility of a relationship, but it is the structural equations that determine whether or not a given relationship is meaningful. The strongest assumptions encoded by a DAG, then, are those edges that are not present. Our goal in this work is to have a language model $\mathcal{P}$ make as many ‘decisions’ about the data generating process as possible. We thus choose this structure as a means of letting $\mathcal{P}$ define whichever structural equations are meaningful or not, and focus on the edge $\tilde{t}\rightarrow\tilde{y}$ as the target for effect estimation. The key criterion we consider for a useful effect estimation benchmark is that the datasets we generate require the use of causal reasoning (e.g., controlling for confounding) in order to recover the effect of $\tilde{t}$ on $\tilde{y}$. Specifically, we aim to generate data for which the observational and interventional distributions are different, i.e., $P^{\mathfrak{B}}_{\tilde{y}\mid\tilde{t}=t}\neq P_{\tilde{y}}^{\mathfrak{B};\text{do}(\tilde{t}=t)}$. In other words, there should be confounders present within $\mathbf{U}\cup\mathbf{X}$. This criterion is not directly in our control given a fixed language model $\mathcal{P}$ [^5] We discuss applications of this same idea to training a model in Section 6. However, even with fixed $\mathcal{P}$ and a fixed DAG structure, we find we are able to achieve it through our choice of sample spaces $\Omega_{\mathbf{U}},\Omega_{\mathbf{X}},\Omega_{\tilde{t}},\Omega_{\tilde{y}}$.
4.1 Breast cancer SD-SCM
We define an SD-SCM over 14 variables in order to explore the effect of a tumor’s PD-L1 expression levels on different breast cancer therapy plans. Rather than reflect medically accurate causal relationships, our goal with this SD-SCM is instead to induce causal structure that can challenge estimation algorithms. Covariates in the breast cancer SD-SCMs are defined in full detail in Appendix A and correspond to the DAG in Figure LABEL:fig:bcancer_dag. For each covariate, 10 different phrasings are considered, resulting in a sample space of $10^{14}$ possible sequences. [^6] Language models are also used to generate the phrasings, but we leave full automation of this process to future work. For example, for the covariate $\tilde{u}_{1}$ that represents ‘age,’ with possible values $\Omega_{\tilde{u}_{1}}=(25,35,45,55,65,75,85)$, two possible phrasings are:
  1. 1.
  2. 2.
We consider 50 different SD-SCM variations, where the sample space for a given SD-SCM is defined by choosing a randomly sampled phrasing from among the possible phrasings for each of the 14 covariates. Then, for each of the 50 SD-SCMs, 20 datasets (each of size 1000) are sampled, for a total of 1000 datasets per language model. In this section, we show results for GPT-2 [49] and Llama-3-8b [17], but we emphasize that the language model is a fully modular component, and thus other language models can be used. Figure 4 shows plots of the features (LABEL:fig:bcancer_hist), propensity scores (LABEL:fig:bcancer_propensity), and ITEs (LABEL:fig:bcancer_ites_binary, LABEL:fig:bcancer_ites) for a single generated dataset using Llama-3-8b. Because the outcome $\tilde{y}$ has $|\Omega_{\tilde{y}}|=4$ possible values, we can consider several possible outcomes, including the observed outcome (categorical), log probabilities for each outcome value (continuous), or probabilities for each outcome value (continuous). This creates, in effect, nine possible targets for each dataset. For benchmarking purposes, we find using probabilities and/or log probabilities as the outcome to be the most useful — there is frequently an effect (even at the individual level) in probability space, even if the sampled outcomes do not change. Comparing Figure LABEL:fig:bcancer_ites_binary to Figure LABEL:fig:bcancer_ites demonstrates this, where for the 400 or so observations where the categorical ITE is zero, the continuous ITE is instead nonzero. Figure LABEL:fig:bcancer_ites also demonstrates that we are able to satisfy our key criterion stated above: the observational distribution $P_{\tilde{y}\mid\tilde{t}=t}^{\mathfrak{B}}$ and interventional distribution $P_{\tilde{y}}^{\mathfrak{B};\text{do}(\tilde{t}=t)}$ are different enough that the SATE and the observed mean difference in outcomes between the treatment and control group are not only different in value, but they also disagree in sign. This is particularly meaningful in a causal inference setting — the treatment appears to lower the outcome, when in fact, its effect is to increase the outcome. We find that how similar $P_{\tilde{y}}^{\mathfrak{B};\text{do}(\tilde{t}=t)}$ is to $P_{\tilde{y}\mid\tilde{t}=t}^{\mathfrak{B}}$ varies across SD-SCMs, explored further by comparing the performance of observational versus casual estimation approaches.

### v2

@@ A.2 Sequence-driven structural causal models
Definition A.6 (Parent-only concatenation).
Given DAG $\mathcal{G}$ over $m$ sequence variables $(\tilde{v}_{1},\ldots,\tilde{v}_{m})$ and a topological ordering $\tau$ consistent with $\mathcal{G}$, parent-only concatenation defines $\left(\tilde{v}_{i}\mid\mathbf{PA}_{i}\right)\equiv\left(\bigoplus_{t\in\mathbf{PA}_{i}}\tilde{v}_{t}\right)\oplus\tilde{v}_{i}$, where $\mathbf{PA}_{i}$ are the parents of $\tilde{v}_{i}$ in $\mathcal{G}$ ordered according to $\tau$.
@@
Definition A.7 (Sequence-driven structural causal model (SD-SCM)).
@@
An SD-SCM uses $\mathcal{P}$ to define an observational distribution over the variables in $\mathbf{V}\cup\mathbf{U}$ that factorizes according to $\mathcal{G}$ via domain-restricted ancestral sampling and parent-only concatenation with $\tau$: $P^{\mathfrak{B}}\equiv\prod_{\tilde{x}_{t}\in\tau}\mathcal{P}_{\tilde{x}_{t}\mid C,\mathbf{PA}_{t}}$. This procedure is shown in LABEL:alg:observational_sampling.
@@
The key difference between an SCM and an SD-SCM is that all variables have at least one common ancestor — the prior inputs $C$ that were used to train the language model, if any. It is however possible to train the LM to induce distributions over the desired variables given this setup. As with the observational distribution, domain-restricted ancestral sampling and parent-only concatenation also allow us to define interventional and counterfactual distributions.
Definition A.8 (Sequence-driven interventional distribution).
@@
An SD-SCM $\mathfrak{B}$ entails the distribution of any subset of variables in $\mathbf{V}\cup\mathbf{U}$ following intervention $I=\text{do}\left(\tilde{v}_{i}=v\right)$ by replacing variable $\tilde{v}_{i}$ with value $v$, and otherwise sampling in the same manner. As with an SCM, we denote an SD-SCM after intervention $I$ as $\mathfrak{B}^{\text{do}(I)}$ and the resulting interventional distribution as $P^{\mathfrak{B};\text{do}(I)}$. This is computed for intervention $\text{do}(\tilde{v}_{i}=v)$ as follows: $P^{\mathfrak{B};\text{do}(\tilde{v}_{i}=v)}\equiv\prod_{\tilde{x}_{t}\in\tau}\mathcal{P}_{\tilde{x}_{t}\mid C,\tilde{v}_{i}=v,\mathbf{PA}^{\prime}_{t}}$, where $\mathbf{PA}^{\prime}_{t}=\mathbf{PA}_{t}\setminus\{\tilde{v}_{i}\}$. This procedure is shown in LABEL:alg:interventional_sampling.
@@
In order to admit unique answers to counterfactual queries, we define abduction for an SD-SCM given evidence $\mathbf{Z}=\mathbf{z}$ as the setting of values $\mathbf{U}=\mathbf{u}$ and any evidence upstream of the intervention, rather than a distribution $P_{\mathbf{U}\mid\mathbf{Z}=\mathbf{z}}$. [^9] Other choices can be explored here, which we leave to future extensions of SD-SCMs. In order to obtain such values $\mathbf{u}$, one needs access to more than just the observed data and language model $\mathcal{P}$ — obtaining $\mathbf{u}$ requires performing bookkeeping during the data generation process. This is a restatement of the fact that computing point counterfactuals in SCMs requires causal mechanisms that are invertible with respect to the noise variables in order to uniquely reconstruct the noise
algorithm]alg:interventional_sampling Inputs: $\text{do}(\tilde{v}_{i}=v),\mathfrak{B}=(\mathbf{V},\mathbf{U},\mathcal{G},\mathcal{P}(\cdot),\tau)$
Returns: $\mathbf{s}=\left(u_{1},\ldots,u_{|\mathbf{U}|},v_{1},\ldots,v_{|\mathbf{V}|}\right)$
$\mathbf{s}\leftarrow()$
for $\tilde{x}_{t}\in\tau$ do
 if $\tilde{x}_{t}\equiv\tilde{v}_{i}$ then
    $x_{t}\leftarrow v$
   end if
 else
    $\textrm{PA}_{\tau}\leftarrow\{t^{\prime}:\tilde{x}_{t^{\prime}}\in\textrm{PA}_{\tilde{x}_{t}}\}$ ordered by $\tau$
    $x_{\textrm{PA}_{\tau}}\leftarrow\bigoplus_{x\in\mathbf{s}\left[\textrm{PA}_{\tau}\right]}x$
    $\mathbf{p}_{\tilde{x}_{t}}\leftarrow[]$
    for $k\in 1,\ldots,\left|\Omega_{\tilde{x}_{t}}\right|$ do
       $x\leftarrow\Omega_{\tilde{x}_{t}}[k]$
       $\mathbf{p}_{\tilde{x}_{t}}[k]\leftarrow\mathcal{P}\left(x_{\textrm{PA}_{\tau}}\oplus x\right)$
      end for
    $P_{\textrm{tot}}\leftarrow\sum_{k}\mathbf{p}_{\tilde{x}_{t}}[k]$
    $j\sim\textrm{Multinomial}(\mathbf{p}_{\tilde{x}_{t}}/P_{\textrm{tot}})$
    $x_{t}\leftarrow\Omega_{\tilde{x}_{t}}[j]$
   end if
 $\textrm{append}(\mathbf{s},x_{t})$
end for
return $\mathbf{s}$
Algorithm 3 A single SD-SCM sample from the interventional distribution for $\text{do}(\tilde{v}_{i}=v)$
that produced a given observation. Because our primary application of SD-SCMs in this work is data generation, such bookkeeping is possible in all our use cases.
Definition A.9 (Sequence-driven counterfactual distribution).
@@
Counterfactual sequence variable $\mathbf{Y}^{*}$ given factual evidence $\mathbf{z}$ and intervention $\text{do}(\mathbf{X}=\mathbf{x})$ (where $\mathbf{X},\mathbf{Y},\mathbf{Z}\subseteq\mathbf{V}$) can be computed for an SD-SCM $\mathfrak{B}$ whenever the exogenous setting $\mathbf{u}=\{u_{1},u_{2},\ldots,u_{|\mathbf{U}|}\}$ that generated evidence $\mathbf{z}$ is known. As with an SCM, for intervention $I$ and observed $\mathbf{Z}=\mathbf{z}$, we denote the counterfactual distribution $P^{\mathfrak{B}\mid\mathbf{Z}=\mathbf{z};\text{do}(I)}$. This is computed for evidence $\mathbf{z}$, exogenous conditions $\mathbf{u}$, and intervention $\text{do}(\tilde{v}_{i}=v)$ as follows: $P^{\mathfrak{B}\mid\mathbf{Z}=\mathbf{z};\text{do}(\tilde{v}_{i}=v)}\equiv\prod_{\tilde{x}_{t}\in\tau}\mathcal{P}_{\tilde{x}_{t}\mid C,\mathbf{U}=\mathbf{u},\mathbf{Z}^{\prime}=\mathbf{z}^{\prime},\tilde{v}_{i}=v,\mathbf{PA}^{\prime\prime}_{t}}$, where $\mathbf{Z}^{\prime}\subseteq\mathbf{Z}$ contains all non-descendants of $\tilde{v}_{i}$ present in $\mathbf{Z}$, and $\mathbf{PA}^{\prime\prime}_{t}=\mathbf{PA}_{t}\setminus\left(\mathbf{U}\cup\mathbf{Z}^{\prime}\cup\{\tilde{v}_{i}\}\right)$. This procedure is shown in LABEL:alg:counterfactual_sampling_appendix.
In plain terms, counterfactual sampling sets not only an intervention $\text{do}(\tilde{v}_{i}=v)$ but also exogenous variables $\mathbf{U}=\mathbf{u}$ and upstream evidence in order to sample a hypothetical alternative that corresponds to the particular unit in question.
In summary, data can be generated from an
algorithm]alg:counterfactual_sampling_appendix Inputs: $\mathbf{s}_{\textrm{obs}}=\left(u_{1},\ldots,u_{|\mathbf{U}|},v_{1},\ldots,v_{|\mathbf{V}|}\right)\newline \text{do}(\tilde{v}_{i}=v),\mathfrak{B}=(\mathbf{V},\mathbf{U},\mathcal{G},\mathcal{P}(\cdot),\tau)$
Returns: $\mathbf{s}^{*}=\left(u_{1},\ldots,u_{|\mathbf{U}|},v^{*}_{1},\ldots,v^{*}_{|\mathbf{V}|}\right)$
$\mathbf{s}^{*}\leftarrow(u_{1},\ldots,u_{|\mathbf{U}|})$
$\textrm{ND}_{i}\leftarrow$ non-descendants of $\tilde{v}_{i}$ in $\mathcal{G}$
for $\tilde{x}_{t}\in\tau\setminus\mathbf{U}$ do
 if $\tilde{x}_{t}\equiv\tilde{v}_{i}$ then
    $x_{t}\leftarrow v$
   end if
 else if $\tilde{x}_{t}\in\textrm{ND}_{i}$ then
    $x_{t}\leftarrow\mathbf{s}_{\textrm{obs}}[t]$
   end if
 else
    $\textrm{PA}_{\tau}\leftarrow\{t^{\prime}:\tilde{x}_{t^{\prime}}\in\textrm{PA}_{\tilde{x}_{t}}\}$ ordered by $\tau$
    $x_{\textrm{PA}_{\tau}}\leftarrow\bigoplus_{x\in\mathbf{s}^{*}\left[\textrm{PA}_{\tau}\right]}x$
    $\mathbf{p}_{\tilde{x}_{t}}\leftarrow[]$
    for $k\in 1,\ldots,\left|\Omega_{\tilde{x}_{t}}\right|$ do
       $x\leftarrow\Omega_{\tilde{x}_{t}}[k]$
       $\mathbf{p}_{\tilde{x}_{t}}[k]\leftarrow\mathcal{P}\left(x_{\textrm{PA}_{\tau}}\oplus x\right)$
      end for
    $P_{\textrm{tot}}\leftarrow\sum_{k}\mathbf{p}_{\tilde{x}_{t}}[k]$
    $j\sim\textrm{Multinomial}(\mathbf{p}_{\tilde{x}_{t}}/P_{\textrm{tot}})$
    $x_{t}\leftarrow\Omega_{\tilde{x}_{t}}[j]$
   end if
 $\textrm{append}(\mathbf{s}^{*},x_{t})$
end for
return $\mathbf{s}^{*}$
Algorithm 4 A single SD-SCM sample from the counterfactual distribution given observation $\mathbf{s}_{\textrm{obs}}$
@@
SD-SCM by domain-restricted forward sampling variables in topological order, and, with adequate bookkeeping, both interventional and counterfactual samples can also be drawn. The key difficulty in generating data this way that is also useful for benchmarking causal inference methods is to ensure it has meaningful structure. In short, it is easy to generate data, but more difficult to generate useful data. The reason for this challenge is that we do not directly specify the structural equations; rather, the structural equations are determined by whatever the language model $\mathcal{P}$ has already encoded.
Appendix B Full description of the breast cancer SD-SCMs